OpenStudy (anonymous):

Two forces with magnitudes of 150 and 100 pounds act on an object at angles of 30° and 120° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

F = 150 (cos(30), sin(30)) + 100 (cos(120), sin(120)) F = 150 (sq3/2, 1/2) + 100 (−1/2, sq3/2) F = (75sq3−50, 75+50sq3) Magnitude of resultant force = sq((75sq3−50)^2 + (75+50sq3)^2) = sq32500 = 50sq13 ≈ 180.278 pounds Direction of resultant force = tan^-1((75+50sq3)/(75sq3−50)) ≈ 63.69 degrees I was just wondering if this was correct...